OpenStudy (moonlight):

simplify 1) In e^2x+10 2) e^2In2x

myininaya (myininaya):

(2) e^(2ln2x)=e^(ln(2x)^2)=(2x)^2=4x^2

OpenStudy (anonymous):

1) 2x + 10 2) e^ln(2x)^2 =(2x)^2=4x^2

OpenStudy (moonlight):

thanx :)